# Drop your Decoder: Pre-training with Bag-of-Word Prediction for Dense Passage Retrieval

### v1

@@ Abstract.
Masked auto-encoder pre-training has emerged as a prevalent technique for initializing and enhancing dense retrieval systems. It generally utilizes additional Transformer decoder blocks to provide sustainable supervision signals and compress contextual information into dense representations. However, the underlying reasons for the effectiveness of such a pre-training technique remain unclear. The usage of additional Transformer-based decoders also incurs significant computational costs. In this study, we aim to shed light on this issue by revealing that masked auto-encoder (MAE) pre-training with enhanced decoding significantly improves the term coverage of input tokens in dense representations, compared to vanilla BERT checkpoints. Building upon this observation, we propose a modification to the traditional MAE by replacing the decoder of a masked auto-encoder with a completely simplified Bag-of-Word prediction task. This modification enables the efficient compression of lexical signals into dense representations through unsupervised pre-training. Remarkably, our proposed method achieves state-of-the-art retrieval performance on several large-scale retrieval benchmarks without requiring any additional parameters, which provides a 67% training speed-up compared to standard masked auto-encoder pre-training with enhanced decoding.  [^1] Our code will be available at https://github.com/ma787639046/bowdpr.
@@
conference: Conference Title; Month xx–xx, xxxx; Conference, Cityccs: Information systems Language models
@@ 1. Introduction
Recent studies have explored several pre-training strategies, such as pre-training with the weakened decoder (Lu et al., 2021), extreme mask ratio (Liu and Shao, 2022), replace token detection (Wang et al., 2022), context reconstruction (Wu et al., 2023), and multi-task decoding (Zhou et al., 2022). As is shown in Figure 1 A), these MAE-style pre-trainings commonly preserve several traits: 1) Asymmetric structure: The encoder is a full-sized BERT for generating discriminative representations based on input sentences. While the decoder is typically a shallow Transformers-based block with a few or just one layer. 2) Bottlenecked connection. The connection between the encoder and the decoder is typically a single representation vector. 3) Reconstruction signals. Carefully designed decoding tasks are applied to the decoder side, which decodes context information through auto-encoding (Gao and Callan, 2021; Liu and Shao, 2022), auto-regression (Lu et al., 2021), enhanced decoding (Liu and Shao, 2022) or other curated decoding tasks (Wang et al., 2022; Zhou et al., 2022; Wu et al., 2022). By optimizing a weak decoder, the lexical information is compressed into the sentence representations, thus providing good initialization ability for retrieval tasks. Notably, enhanced decoding from (Liu and Shao, 2022) is proven effective for dense retrieval pre-training, which utilizes sentence representations as query streams and masked context embeddings as content streams.
However, we argue that existing MAE-style pre-training paradigms have several drawbacks, which may hinder the development of dense retrieval pre-training: 1) Lack of interpretability. Existing studies focus on creating carefully designed tasks, but still neglect to reveal the underlying reason for why such a pre-training method is effective. 2) Additional complexity. MAE-style pre-training involves optimization with millions of document collections. The usage of additional Transformers-based decoders brings considerable GPU memory usage and additional $O(n^{2})$ computational complexity. Some studies also need additional data preprocessing (Wu et al., 2022; Zhou et al., 2022) or extensive data masking (Liu and Shao, 2022; Wu et al., 2023).
In this study, we aim to solve the above issues and develop a more computationally efficient and more interpretable pre-training technique. The semantic meaning of a dense representation is typically condensed in latent spaces and hard to directly interpret. In our preliminary studies, we project the low-dimensional representation to the vocabulary space with the BERT language model heads and compare the changes in MAE-style pre-training. Notably, we discover that MAE-style pre-training with enhanced decoding (Liu and Shao, 2022) significantly enhances the term coverage of original input tokens, which implies that the stronger retrieval ability may come from the compression of input lexicon terms into dense representations.
@@
$\bullet$ High interpretability. Compared to complex modifications to the decoder tasks (Liu and Shao, 2022; Wang et al., 2022; Wu et al., 2023; Zhou et al., 2022) of MAE-style pre-training, our method removes the need for adding decoders and directly optimizes a multi-label prediction task of input lexicons.
To verify the effectiveness of our method, we conduct experiments on several large-scale web search and open-domain QA datasets, including MS-MARCO passage ranking (Nguyen et al., 2016), Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017) and heterogeneous zero-shot BEIR benchmarks (Thakur et al., 2021). Our proposed method achieves state-of-the-art retrieval performance and provides a 67% training speed-up compared to masked auto-encoder with enhanced decoding (Liu and Shao, 2022).
@@ 2. Related Works
Passage retrieval aims to find relevant information from large passage collections based on a given search query. Traditional techniques for ad-hoc retrieval, like TF-IDF and BM25 (Robertson et al., 2009), perform relevant searches based on lexicon terms and inverted document frequencies. Such methods have efficient retrieval abilities, but they suffer from vocabulary mismatch problems (Karpukhin et al., 2020). Although there are continuous techniques like query expansions (Nogueira et al., 2019; Wang et al., 2023; Jagerman et al., 2023) and query rewrite (Gollapudi et al., 2011) to alleviate such mismatch issues, additional preprocess of queries or passage collections make the retrieval pipelines much more complicated and hard to maintain.
@@
Dense retrieval (Karpukhin et al., 2020; Qu et al., 2021) has become popular in mainstream retrieval applications recently, because it maps the textual information into learnable representations and naturally alleviates the lexicon mismatch problems by learning the similarity traits of queries and passages through end-to-end contrastive learning. Dense retrieval encodes queries and passage collections as condensed low-dimension sentence representations with pre-trained language models (PLMs), and performs retrieval based on the similarity of representations with efficient search algorithms like maximum inner-product search (MIPS) (Mussmann and Ermon, 2016).
Pre-training for Dense Retrieval.
There are continuous efforts (Gao and Callan, 2021, 2022; Ren et al., 2021; Lu et al., 2022) to improve the retrieval quality of sentence representations with specifically designed pre-training techniques, where masked auto-encoder (MAE) (Gao and Callan, 2021; Lu et al., 2021; Gao and Callan, 2022; Liu and Shao, 2022; Wu et al., 2023; Wang et al., 2022; Zhou et al., 2022) has gained popularity to boost the initialization of PLMs for downstream retrieval tasks in recent years. Masked auto-encoder (MAE) pre-training seeks to compress the lexicon information of contextual texts into sentence representations through various auxiliary pre-training tasks and tailored bottlenecked architectures. It usually appends a shallow decoder to the PLM encoder and connects the encoder-decoder with merely a sentence representation. Through the process of MAE-style pre-training, the information of input texts is first condensed as sentence representations and then decoded in the appended decoder with auto-encoding (Gao and Callan, 2021, 2022; Liu and Shao, 2022; Wu et al., 2023; Zhou et al., 2022), auto-regression (Lu et al., 2021) or enhanced decoding (Liu and Shao, 2022). Such an information encoding-decoding process brings sustainable supervision signals to the sentence representations and thus gives better initialization to the PLMs for downstream retrievals.
Recent studies focus on designing various decoding tasks for MAE-based pre-trainings. (Lu et al., 2021) proposes to pre-train an encoder with merely a single-layer auto-regressive decoder (Radford et al., 2019) and apply attention window restriction to enforce textual information compression to the representations. (Gao and Callan, 2021, 2022) divides the MAE into early-late-head layers and proposes to optimize with condensed representations and structural readiness. Furthermore, (Liu and Shao, 2022; Liu et al., 2023) proposes pre-training with an auto-encoding decoder and an extreme mask ratio. They further propose an enhanced decoding paradigm (Yang et al., 2019) for improving token usage ability in MAE-style pre-training. It applies sentence representations with positional information as query streams and performs cross-attention with masked context embeddings as content streams in the decoder. In our studies below, we observe significantly increased term coverage of input lexicon tokens by projecting the pre-trained sentence representations of enhanced decoding to the vocabulary space. This implies that the stronger retrieval ability may come from the mediate compression of lexicon terms into dense representations.
Although the bloom of MAE-style pre-training gives better retrieval performances to dense systems, several obvious drawbacks hinder the further development of such pre-training techniques. Firstly, existing MAE pre-trainings focus on creating carefully designed auxiliary pre-training tasks, but it’s hard to interpret the underlying reason why such an auxiliary pre-training task is effective. Furthermore, MAE-style pre-training mandatory needs the additional Transformers-based decoders, which brings considerable GPU memory cost and additional $O(n^{2})$ computational complexity. Complex data preprocessing (Wu et al., 2022; Zhou et al., 2022) or extensive data masking (Liu and Shao, 2022; Wu et al., 2023) are also needed in several implementations. In contrast, our method breaks the constraints of using any additional decoders, preprocessing, or masking, by directly optimizing a Bag-of-Word decoding task. Such a method achieves state-of-the-art retrieval performances on various datasets. It is highly interpretable and has nearly zero additional computational complexity in the whole pre-training process.
@@ 3.1. Preliminary of Dense Retrieval Pre-training
Decoder designs are key innovations of MAE-style pre-training. There are mainly three categories of decoder designs, including auto-encoding (Gao and Callan, 2021; Liu and Shao, 2022), auto-regression (Lu et al., 2021), and enhanced decoding (Liu and Shao, 2022).
@@
Figure 2. Compositions of Top-k tokens of dense representation.
@@
Unlike self-attention schemas with auto-encoding or auto-regression, enhanced decoding (Liu and Shao, 2022; Yang et al., 2019) develops a cross-attention pattern with an extreme masked ratio for improving token usage. The decoder inputs $t_{Dec}$ are first forwarded through the embedding layer and converted to input embeddings $h_{Dec}^{Emb}$. Then enhanced decoding utilizes encoder representations $h_{last}^{CLS}$ with positional embeddings $p$ as the query stream $q$, and input embeddings $h_{Dec}^{Emb}$ as the content stream $k$ and $v$.
@@
Specifically, a position-specific attention mask $M_{ij}\in[L,L]$ ($L$ is the sequence length) is applied to the cross-attention to increase the difficulty of the decoder optimization objective. This mask $M_{ij}$ is composed of masking diagnose positions $i=j$ and a randomly sampled token mask $i,j\in sample([0...L])$. After projecting the decoder’s hidden states $h_{Dec}$ with BERT language-model heads (i.e., its embedding matrix $E^{T}$), a cross-entropy loss is used with all corresponding input tokens as labels.
@@
Algorithm 2 Bag-of-word Prediction
0:  Tokenized Inputs $T$
2:  $h_{proj}\leftarrow h_{last}^{CLS}\cdot E^{T}$triangleright Projection
3:  Bag-of-word Prediction Loss $\mathcal{L}_{Dec}$
3.2. Interprete MAE Pre-training
Adaption Abilities and Retrieval Performances
The MLM loss on target corpora reflects the modeling abilities of various pre-trained checkpoints. To evaluate the domain adaption abilities of the PLMs on the target corpus, we randomly sample 10,000 corpora from MS-MARCO (Nguyen et al., 2016) and DPR-Wiki (Karpukhin et al., 2020) collections and calculate the corresponding average MLM loss.
Here we involve 5 models for performance comparison. 1) Original BERT checkpoint (Devlin et al., 2019); 2) The BERT further pre-trained with MLM on the original training corpus, i.e., Wikipedia and Bookcorpus collections; 3) MAE Pre-training with Auto-Encoding Decoder; 4) MAE Pre-training with Auto-Regression Decoder; 5) MAE Pre-training with Enhanced Decoding Decoder. For fair comparisons, we mainly reproduce the pre-training settings in (Liu and Shao, 2022) and directly take the enhanced decoding checkpoint from (Liu and Shao, 2022). The MLM loss on evaluation sets of MS-MARCO, DPR-Wiki, and the corresponding retrieval performances are listed in Table 1 and 2.
Here we can observe some interesting phenomena. 1) Both further pre-training with pure MLM loss and MAE-style pre-training significantly lower the MLM loss on the target evaluation corpus, and improve the downstream retrieval performances. 2) Although MAE-style pre-training has higher MLM loss than the further pre-trained checkpoint, they tend to achieve higher retrieval performances. Thus we can conclude that pre-training for dense retrieval is effective for better domain adaption abilities and retrieval performances, but such abilities are not always positively correlated to lower MLM losses on target corpora when using MAE-style pre-training. Thus, other critical traits should contribute to better retrieval abilities for MAE-style pre-training.

### v2

@@ Abstract.
Masked auto-encoder pre-training has emerged as a prevalent technique for initializing and enhancing dense retrieval systems. It generally utilizes additional Transformer decoder blocks to provide sustainable supervision signals and compress contextual information into dense representations. However, the underlying reasons for the effectiveness of such a pre-training technique remain unclear. The usage of additional Transformer-based decoders also incurs significant computational costs. In this study, we aim to shed light on this issue by revealing that masked auto-encoder (MAE) pre-training with enhanced decoding significantly improves the term coverage of input tokens in dense representations, compared to vanilla BERT checkpoints. Building upon this observation, we propose a modification to the traditional MAE by replacing the decoder of a masked auto-encoder with a completely simplified Bag-of-Word prediction task. This modification enables the efficient compression of lexical signals into dense representations through unsupervised pre-training. Remarkably, our proposed method achieves state-of-the-art retrieval performance on several large-scale retrieval benchmarks without requiring any additional parameters, which provides a 67% training speed-up compared to standard masked auto-encoder pre-training with enhanced decoding.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657792isbn: 979-8-4007-0431-4/24/07ccs: Information systems Language models
@@ 1. Introduction
Recent studies have explored several pre-training strategies, such as pre-training with the weakened decoder (Lu et al., 2021), extreme mask ratio (Xiao et al., 2022), replace token detection (Wang et al., 2023b), context reconstruction (Wu et al., 2023a), and multi-task decoding (Zhou et al., 2022). As is shown in Figure 1 A), these MAE-style pre-trainings commonly preserve several traits: 1) Asymmetric structure: The encoder is a full-sized Transformers-based encoder for generating discriminative representations based on input sentences. While the decoder is typically a shallow Transformers-based block with a few or just one layer. 2) Bottlenecked connection. The connection between the encoder and the decoder is typically a single representation vector. 3) Reconstruction signals. Carefully designed decoding tasks are applied to the decoder side, which decodes context information through auto-encoding (Gao and Callan, 2021; Xiao et al., 2022), auto-regression (Lu et al., 2021), enhanced decoding (Xiao et al., 2022) or other curated decoding tasks (Wang et al., 2023b; Zhou et al., 2022; Wu et al., 2023b). By optimizing a weak decoder, the lexical information is compressed into the sentence representations, thus providing good initialization ability for retrieval tasks. Notably, enhanced decoding from (Xiao et al., 2022) is proven effective for dense retrieval pre-training, which utilizes sentence representations as query streams and masked context embeddings as content streams.
However, we argue that existing MAE-style pre-training paradigms have several drawbacks, which may hinder the development of dense retrieval pre-training: 1) Lack of interpretability. Existing studies focus on creating carefully designed tasks, but still neglect to reveal the underlying reason for why such a pre-training method is effective. 2) Additional complexity. MAE-style pre-training involves optimization with millions of document collections. The usage of additional Transformers-based decoders brings considerable GPU memory usage and additional $O(n^{2})$ computational complexity. Some studies also need additional data preprocessing (Wu et al., 2023b; Zhou et al., 2022) or extensive data masking (Xiao et al., 2022; Wu et al., 2023a).
In this study, we aim to solve the above issues and develop a more computationally efficient and more interpretable pre-training technique. The semantic meaning of a dense representation is typically condensed in latent spaces and hard to directly interpret. In our preliminary studies, we project the low-dimensional representation to the vocabulary space with the language model heads and compare the changes in MAE-style pre-training. Notably, we discover that MAE-style pre-training with enhanced decoding (Xiao et al., 2022) significantly enhances the term coverage of original input tokens, which implies that the stronger retrieval ability may come from the compression of input lexicon terms into dense representations.
@@
$\bullet$ High interpretability. Compared to complex modifications to the decoder tasks (Xiao et al., 2022; Wang et al., 2023b; Wu et al., 2023a; Zhou et al., 2022) of MAE-style pre-training, our method removes the need for adding decoders and directly optimizes a multi-label prediction task of input lexicons.
To verify the effectiveness of our method, we conduct experiments on several large-scale web search and open-domain QA datasets, including MS-MARCO passage ranking (Nguyen et al., 2016), Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017) and heterogeneous zero-shot BEIR benchmarks (Thakur et al., 2021). Our proposed method achieves state-of-the-art retrieval performance and provides a 67% training speed-up compared to masked auto-encoder with enhanced decoding (Xiao et al., 2022).
@@ 2. Related Works
Passage retrieval aims to find relevant information from large passage collections based on a given search query. Traditional techniques for ad-hoc retrieval, such as TF-IDF and BM25 (Robertson et al., 2009), perform relevant searches based on lexicon terms and inverted document frequencies. Such methods have efficient retrieval abilities, but they suffer from vocabulary mismatch problems (Karpukhin et al., 2020). Although there are continuous techniques like query expansions (Nogueira et al., 2019; Wang et al., 2023a; Jagerman et al., 2023) and query rewrite (Gollapudi et al., 2011) to alleviate such mismatch issues, additional preprocess of queries or passage collections make the retrieval pipelines much more complicated and hard to maintain.
@@
Dense retrieval (Karpukhin et al., 2020; Qu et al., 2021) has become popular in mainstream retrieval applications recently, because it maps the textual information into learnable representations and naturally alleviates the lexicon mismatch problems by learning the similarity traits of queries and passages through end-to-end contrastive learning. Dense retrieval encodes queries and passage collections as condensed low-dimension sentence representations with pre-trained language models (PLMs), and performs retrieval based on the similarity of representations with efficient search algorithms like maximum inner-product search (MIPS) (Mussmann and Ermon, 2016). Notably, there are continuous efforts (Formal et al., 2021b, a; Shen et al., 2023; Yu et al., 2024) to adapt the PLMs to sparse retrieval. There are also extended studies about hybrid retrievals (Liu et al., 2023; Wu et al., 2023c), which combines both dense and sparse retrievals to bring superior retrieval performances. However, our work primarily focuses on dense retrieval pre-training and is orthogonal to these works.
@@
Pre-training for Dense Retrieval.
There are continuous efforts (Gao and Callan, 2021, 2022; Ren et al., 2021; Lu et al., 2022) to improve the retrieval quality of sentence representations with specifically designed pre-training techniques, where masked auto-encoder (MAE) (Gao and Callan, 2021; Lu et al., 2021; Gao and Callan, 2022; Xiao et al., 2022; Wu et al., 2023a; Wang et al., 2023b; Zhou et al., 2022) has gained popularity to boost the initialization of PLMs for downstream retrieval tasks in recent years. Masked auto-encoder (MAE) pre-training seeks to compress the lexicon information of contextual texts into sentence representations through various auxiliary pre-training tasks and tailored bottlenecked architectures. It usually appends a shallow decoder to the PLM encoder and connects the encoder-decoder with merely a sentence representation. Through the process of MAE-style pre-training, the information of input texts is first condensed as sentence representations and then decoded in the appended decoder with auto-encoding (Gao and Callan, 2021, 2022; Xiao et al., 2022; Wu et al., 2023a; Zhou et al., 2022), auto-regression (Lu et al., 2021) or enhanced decoding (Xiao et al., 2022). Such an information encoding-decoding process brings sustainable supervision signals to the sentence representations and thus gives better initialization to the PLMs for downstream retrievals.
Recent studies focus on designing various decoding tasks for MAE-based pre-trainings. (Lu et al., 2021) proposes to pre-train an encoder with merely a single-layer auto-regressive decoder (Radford et al., 2019) and apply attention window restriction to enforce textual information compression to the representations. (Gao and Callan, 2021, 2022) divides the MAE into early-late-head layers and proposes to optimize with condensed representations and structural readiness. Furthermore, (Xiao et al., 2022; Liu et al., 2023) proposes pre-training with an auto-encoding decoder and an extreme mask ratio. They further propose an enhanced decoding paradigm (Yang et al., 2019) for improving token usage ability in MAE-style pre-training. It applies sentence representations with positional information as query streams and performs cross-attention with masked context embeddings as content streams in the decoder. In our studies below, we observe significantly increased term coverage of input lexicon tokens by projecting the pre-trained sentence representations of enhanced decoding to the vocabulary space. This implies that the stronger retrieval ability may be attributed to the compression of lexicon terms into dense representations.
@@
Drawbacks of Existing MAE Pre-training.
Although the bloom of MAE-style pre-training gives better retrieval performances to dense systems, several obvious drawbacks hinder the further development of such pre-training techniques. Firstly, existing MAE pre-trainings focus on creating carefully designed auxiliary pre-training tasks, but it’s hard to interpret the underlying reason why such an auxiliary pre-training task is effective. Furthermore, MAE-style pre-training mandatory needs the additional Transformers-based decoders, which brings considerable GPU memory cost and additional $O(n^{2})$ computational complexity. Complex data preprocessing (Wu et al., 2023b; Zhou et al., 2022) or extensive data masking (Xiao et al., 2022; Wu et al., 2023a) are also needed in several implementations. In contrast, our method breaks the constraints of using any additional decoders, preprocessing, or masking, by directly optimizing a Bag-of-Word decoding task. Such a method achieves state-of-the-art retrieval performances on various datasets. It is highly interpretable and has nearly zero additional computational complexity in the whole pre-training process. Note that some works also involve term prediction tasks (Zhuang and Zuccon, 2021; Déjean et al., 2023). However, (Zhuang and Zuccon, 2021) utilizes the query-document likelihoods in the fine-tuning stage, while our work utilizes a simpler Bag-of-Word prediction task without query-document interaction. Our study also focuses on the interpretability of MAE pre-training for dense retrieval and optimization in unsupervised pre-training. (Déjean et al., 2023) makes performance comparisons with recent middle training methods and augments with the prediction of L2 normalized frequencies of the input tokens. In comparison, our work reveals the underlying mechanism of performance gains for MAE pre-training and proposes a simple Bag-of-Word prediction instead of frequency prediction.
@@ 3.1. Preliminary of Dense Retrieval Pre-training
Decoder designs are key innovations of MAE-style pre-training. There are mainly three categories of decoder designs, including auto-encoding (Gao and Callan, 2021; Xiao et al., 2022), auto-regression (Lu et al., 2021), and enhanced decoding (Xiao et al., 2022).
@@
Algorithm 2 Bag-of-word Prediction
0:  Tokenized Inputs $T$
2:  $h_{proj}\leftarrow h_{last}^{CLS}\cdot E^{T}$triangleright Projection
3:  Bag-of-word Prediction Loss $\mathcal{L}_{Dec}$
@@
Unlike self-attention schemas with auto-encoding or auto-regression, enhanced decoding (Xiao et al., 2022; Yang et al., 2019) develops a cross-attention pattern with an extreme masked ratio for improving token usage. The decoder inputs $t_{Dec}$ are first forwarded through the embedding layer and converted to input embeddings $h_{Dec}^{Emb}$. Then enhanced decoding utilizes encoder representations $h_{last}^{CLS}$ with positional embeddings $p$ as the query stream $q$, and input embeddings $h_{Dec}^{Emb}$ as the content stream $k$ and $v$.
@@
Specifically, a position-specific attention mask $M_{ij}\in[L,L]$ ($L$ is the sequence length) is applied to the cross-attention to increase the difficulty of the decoder optimization objective. This mask $M_{ij}$ is composed of a diagnose positional mask $i=j$ and a randomly sampled token mask $i,j\in sample([0...L])$. After projecting the decoder’s hidden states $h_{Dec}$ with BERT language-model heads (i.e., its embedding matrix $E^{T}$), a cross-entropy loss is used with all corresponding input tokens as labels.
@@
3.2. Interpret MAE Pre-training
Adaptation Abilities and Retrieval Performances
The MLM loss on target corpora reflects the modeling abilities of various pre-trained checkpoints. To evaluate the domain adaptation abilities of the PLMs on the target corpus, we randomly sample 10,000 passages from MS-MARCO (Nguyen et al., 2016) and DPR-Wiki (Karpukhin et al., 2020) collections and calculate the corresponding average MLM loss.
Here we involve 5 models for performance comparison. 1) Original BERT checkpoint (Devlin et al., 2019); 2) The BERT further pre-trained with MLM on the original training corpus, i.e., Wikipedia and Bookcorpus collections; 3) MAE Pre-training with Auto-Encoding Decoder; 4) MAE Pre-training with Auto-Regression Decoder; 5) MAE Pre-training with Enhanced Decoding Decoder. For fair comparisons, we mainly reproduce the pre-training settings in (Xiao et al., 2022) and directly take the enhanced decoding checkpoint from it. The MLM loss on evaluation sets of MS-MARCO, DPR-Wiki, and the corresponding retrieval performances are listed in Table 1 and 2.
Here we can observe some interesting phenomena. 1) Both further pre-training with pure MLM loss and MAE-style pre-training significantly lower the MLM loss on the target evaluation corpus, and improve the downstream retrieval performances. 2) Although MAE-style pre-training has higher MLM loss than the further pre-trained checkpoint, they tend to achieve higher retrieval performances. Thus we can conclude that pre-training for dense retrieval is effective for better domain adaptation abilities and retrieval performances, but such abilities are not always positively correlated to lower MLM losses on target corpora when using MAE-style pre-training. Thus, other critical traits should contribute to better retrieval abilities for MAE-style pre-training.
Additional MS-MARCO Natural Question Trivia QA
Model Decoder MRR@10 R@50 R@1k R@5 R@20 R@100 R@5 R@20 R@100
BM25 (Robertson et al., 2009) 18.7 59.2 85.7 - 59.1 73.7 - 66.9 76.7
SPLADE++ (Formal et al., 2022) 38.0 - 98.2 - - - - - -
DPR (Karpukhin et al., 2020) 31.7 58.0 85.7 - 74.4 85.3 - 79.3 84.9
ANCE (Xiong et al., 2021) 33.0 - 95.9 - 81.9 87.5 - 80.3 85.3
SEED (Lu et al., 2021) $\checkmark$ 33.9 - 96.1 - 83.1 88.7 - - -
RocketQA (Qu et al., 2021) 37.0 - 97.9 74.0 82.7 88.5 - - -
SimLM (Wang et al., 2023b) $\checkmark$ 39.1 87.3 98.6 73.6 84.3 89.0 78.1 83.8 87.6
RetroMAE (Xiao et al., 2022) $\checkmark$ 39.3 87.0 98.5 74.4 84.4 89.4 78.9 84.5 88.0
BoW Prediction 40.1* 88.7* 98.9 75.3 84.6 90.4* 79.4* 84.9* 88.0
Data Process Additional Decoder Training Speed
Further Pre-train Encoder $O(n)$ 0.0476 - - 269.708 -
Auto-Encoding (Gao and Callan, 2021, 2022; Xiao et al., 2022; Wu et al., 2023a) Encoder-Decoder $O(n)$ 0.0940 $O(n^{2})$ 0.0013 222.658 17.4%
Enhanced Decoding (Xiao et al., 2022) Encoder-Decoder $O(n^{2})$ 5.6261 $O(n^{2})$ 0.0012 85.797 68.2%
BoW Prediction Encoder $O(n)$ 0.0533 - 0.0002 266.359 1.2%
Table 4. Comparisons of architecture, data process, additional decoder, and training speed of different pre-training methods. All measurements are conducted on the same environment with a single NVIDIA H800 GPU. Note that Bag-of-Word (BoW) prediction does not involve additional decoders, thus we report the GPU computation time of the prediction task. The original implementation of enhanced decoding (Xiao et al., 2022) randomly generates 256 attention masks for each input text, and picks one as the final mask used in the pre-training. And thus it has $O(n^{2})$ data process complexity.